# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v1

@@ Abstract
We show the phenomenon of infinite coherence cost by obtaining the general lower bounds for the coherence cost required to approximately implement a certain class of Gibbs-preserving Operations, which diverges at the limit of zero implementation error. We show that our lower bound is almost tight, where we in turn find that a quantifier introduced in Ref. Tajima et al. (2022), which is linked to the capability of changing energy, characterizes the optimal coherence cost for certain Gibbs-preserving Operations. We also find that an arbitrary (not necessarily Gibbs-preserving) quantum channel can generally be approximately implemented by Thermal Operations with a coherence cost that scales with the error in the same way as the aforementioned lower bound for Gibbs-preserving Operations, showing that some Gibbs-preserving Operations are, roughly speaking, belong to the most costly class of quantum operation.
@@
It is insightful to see the structural difference between the Gibbs-preserving Operations in Eqs. (2) and (10). The one in (2) can create coherence from an incoherent state input state $|{1}\rangle\!\langle{1}|$. On the other hand, the channel in (10) cannot create coherence at all—in fact, output states are always incoherent for any input states. Instead, it can perform a measurement in a coherent basis. As we show in Appendix D, the coherent cost for the channel in (2) is upper bounded by ${\cal F}(\eta)+{\cal F}(\sigma)$, which corresponds to the sum of coherence that can be created by the channel. This shows a drastic asymmetry between creation and detection of coherence when it comes to its realization.
@@
Proof can be found in Appendix E. This particularly ensures that the quantity ${\cal C}(\Lambda,\mathbb{P})$ defined in (6) serves as a key quantity that characterizes the coherence cost for a certain class of Gibbs-preserving Operations, which in turn provides an operational interpretation to this quantity.
Theorem 4 ensures the existence of a Gibbs-preserving Operation that almost achieves the lower bound. However, it does not tell much about the general upper bound that could be applied to an arbitrary Gibbs-preserving Operation. In the following, we show that, by giving up obtaining the form that almost matches the lower bound, we can obtain the general sufficient coherence cost that can be universally applied to all Gibbs-preserving Operations. In fact, we find that the applicability of our bound is much beyond Gibbs-preserving Operations—it gives a sufficient coherence cost for an arbitrary quantum channel. (Proof in Appendix E.)
@@
Let us now remark a unique characteristic of our results in relation to the previous result for trivial Hamiltonian. When input and output states are only equipped with trivial Hamiltonian, every state becomes an incoherent state, i.e., invariant under Hamiltonian evolution, and thus coherence loses its status as a precious resource. Therefore, the meaningful question in such a setting is to ask the work cost (the minimum number of work bit required) for implementing unital channels using Noisy Operations Horodecki et al. (2003), which respectively corresponds to the Gibbs-preserving and Thermal Operations for trivial Hamiltonian. Ref. Faist et al. (2015b) showed that the minimum work cost is upper bounded by a quantity scaling as $\sim\log(1/\epsilon)$ with implementation error $\epsilon$, which particularly diverges at the limit of exact implementation. Nevertheless, its lower bound has still not been established, and therefore it is still unclear if this diverging cost is a fundamental phenomenon or merely an artifact of their specific construction, which is based on a decoupling technique Rio et al. (2011). On the other hand, our Theorem 2 provides a lower bound that scale with $1/\epsilon$, which is complemented by upper bounds in Theorems 4 and 5 with the same scaling. To the best of our knowledge, our results are the first ones that establish the inherently diverging thermodynamic cost for implementing Gibbs-preserving Operations.
@@
Acknowledgments.— We thank Kaito Watanabe for discussions. H.T. was supported by JSPS Grants-in-Aid for Scientific Research No. JP19K14610, No. JP22H05250, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061. R.T. acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JP24K16975, JST, CREST Grant Number JPMJCR23I3, Japan, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” Grant Number JP24H00943.
@@ A.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
We remark that the above relation can be extended to a general state ensemble $\{p_{j},\rho_{j}\}_{j}$ that may not be orthogonal to each other Tajima et al. (2022). The relation (31) is given as a unification between the Wigner-Araki-Yanase (WAY) theorems on quantum measurements Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023) and unitary gates Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021). It also allows restrictions on the classical information recovery in the Hayden-Preskill thought experiments Hayden and Preskill (2007) imposed by the energy conservation Tajima et al. (2022) and extends the WAY theorem to various errors and disturbances of quantum measurements and the out-of-time-ordered correlators Emori and Tajima (2023). Here, we utilize this relation to obtain a lower bound for coherence cost for approximately implementing Gibbs-preserving Operations by Thermal Operations.
@@ Appendix C Construction of pairwise reversible Gibbs-preserving Operations
For a countably infinite series $\{E_{n}\}_{n}$ of real numbers, let $S_{d}(\{E_{n}\}_{n})$ be an arbitrary $d$-dimensional system equipped with Hamiltonian $H_{d}=\sum_{n=0}^{d-1}E_{n}|{n}\rangle\!\langle{n}|$. Then, for arbitrary $d\geq 3$ and $d^{\prime}\leq d$, and an arbitrary energy spectrum $\{E_{n}\}_{n}$ with $E_{i}\geq E_{j},\ \forall i,j$ such that it is not fully degenerate above the ground energy, i.e., there exists $11\leq i\leq d-$ 1 such that $E_{i+1}>E_{i}$, there exists a pairwise reversible Gibbs-preserving map $\Lambda:S_{d}(\{E_{n}\}_{n})\to S_{d^{\prime}}(\{E_{n}\}_{n})$ with a reversible pair $\mathbb{P}$ such that ${\cal C}(\Lambda,\mathbb{P})>0$.
@@
Appendix D Upper bound for the coherence cost of Eq. (2)

### v2

@@ Abstract
We show the phenomenon of infinite coherence cost by obtaining the general lower bounds for the coherence cost required to approximately implement a certain class of Gibbs-preserving Operations, which diverges at the limit of zero implementation error. We show that our lower bound is almost tight, where we in turn find that a quantifier introduced in Ref. Tajima et al. (2022), which is linked to the capability of changing energy, characterizes the optimal coherence cost for certain Gibbs-preserving Operations. We also find that an arbitrary (not necessarily Gibbs-preserving) quantum channel can generally be approximately implemented by Thermal Operations with a coherence cost that scales with the error in the same way as the aforementioned lower bound for Gibbs-preserving Operations, showing that some Gibbs-preserving Operations, roughly speaking, belong to the most costly class of quantum operation.
@@
It is insightful to see the structural difference between the Gibbs-preserving Operations in Eqs. (2) and (10). The one in (2) can create coherence from an incoherent state input state $|{1}\rangle\!\langle{1}|$. On the other hand, the channel in (10) cannot create coherence at all—in fact, output states are always incoherent for any input states. Instead, it can perform a measurement in a coherent basis. As we show in Appendix E, the coherent cost for the channel in (2) is upper bounded by ${\cal F}(\eta)+{\cal F}(\sigma)$, which corresponds to the sum of coherence that can be created by the channel. This shows a drastic asymmetry between creation and detection of coherence when it comes to its realization.
@@
We furthermore find that our findings restrict the physically feasible state transformations. The following result shows that whenever the systems $S$ and $S^{\prime}$ satisfy the same condition as that for Theorem 3, there exist state transformations that can be achieved by Gibbs-preserving Operations but none of those Gibbs-preserving Operations cannot be implemented by a finite amount of coherence together with Thermal Operations.
Theorem 4.
Suppose the input and output systems $S$ and $S^{\prime}$ satisfy the condition (9). Then, there exists a pair $(\rho,\sigma)$ of states such that the state transition $\rho\rightarrow\sigma$ is possible by a Gibbs-preserving Operation but every Gibbs-preserving Operation $\Lambda$ such that $\Lambda(\rho)=\sigma$ must come with diverging coherence cost, i.e., ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$.
We prove this in Appendix D, which also provides an explicit expression of the pair $(\rho,\sigma)$ of states. Theorem 4 shows that when the available coherence resource is restricted to be finite, the range of transformations achievable by Gibbs-preserving Operations is strictly narrowed. We stress that the condition (9) is extremely mild—for virtually all scenarios Theorem 4 puts nontrivial restrictions on the physically available state transformations by Gibbs-preserving Operations.
@@
Proof can be found in Appendix F. This particularly ensures that the quantity ${\cal C}(\Lambda,\mathbb{P})$ defined in (6) serves as a key quantity that characterizes the coherence cost for a certain class of Gibbs-preserving Operations, which in turn provides an operational interpretation to this quantity.
Theorem 5 ensures the existence of a Gibbs-preserving Operation that almost achieves the lower bound. However, it does not tell much about the general upper bound that could be applied to an arbitrary Gibbs-preserving Operation. In the following, we show that, by giving up obtaining the form that almost matches the lower bound, we can obtain the general sufficient coherence cost that can be universally applied to all Gibbs-preserving Operations. In fact, we find that the applicability of our bound is much beyond Gibbs-preserving Operations—it gives a sufficient coherence cost for an arbitrary quantum channel. (Proof in Appendix F.)
@@
Let us now remark a unique characteristic of our results in relation to the previous result for trivial Hamiltonian. When input and output states are only equipped with trivial Hamiltonian, every state becomes an incoherent state, i.e., invariant under Hamiltonian evolution, and thus coherence loses its status as a precious resource. Therefore, the meaningful question in such a setting is to ask the work cost (the minimum number of work bit required) for implementing unital channels using Noisy Operations Horodecki et al. (2003), which respectively corresponds to the Gibbs-preserving and Thermal Operations for trivial Hamiltonian. Ref. Faist et al. (2015b) showed that the minimum work cost is upper bounded by a quantity scaling as $\sim\log(1/\epsilon)$ with implementation error $\epsilon$, which particularly diverges at the limit of exact implementation. Nevertheless, its lower bound has still not been established, and therefore it is still unclear if this diverging cost is a fundamental phenomenon or merely an artifact of their specific construction, which is based on a decoupling technique Rio et al. (2011). On the other hand, our Theorem 2 provides a lower bound that scale with $1/\epsilon$, which is complemented by upper bounds in Theorems 5 and 6 with the same scaling. To the best of our knowledge, our results are the first ones that establish the inherently diverging thermodynamic cost for implementing Gibbs-preserving Operations.
@@
Acknowledgments.— We thank Kaito Watanabe and Bartosz Regula for fruitful discussions. H.T. was supported by JSPS Grants-in-Aid for Scientific Research No. JP19K14610 and No. JP22H05250, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas B “Quantum Energy Innovation” Grant Numbers 24H00830 and 24H00831, and JST PRESTO No. JPMJPR2014, JST MOONSHOT No. JPMJMS2061. R.T. acknowledges the support of JSPS KAKENHI Grant Number JP23K19028, JP24K16975, JST, CREST Grant Number JPMJCR23I3, Japan, and MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” Grant Number JP24H00943.
@@ A.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
We remark that the above relation can be extended from $\mathbb{P}$ to a general state ensemble $\Omega:=\{p_{j},\rho_{j}\}_{j}$ that may not be orthogonal to each other Tajima et al. (2022). The relation (31) is given as a unification between the Wigner-Araki-Yanase (WAY) theorems on quantum measurements Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023) and unitary gates Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021). It also allows restrictions on the classical information recovery in the Hayden-Preskill thought experiments Hayden and Preskill (2007) imposed by the energy conservation Tajima et al. (2022) and extends the WAY theorem to various errors and disturbances of quantum measurements and the out-of-time-ordered correlators Emori and Tajima (2023). Here, we utilize this relation to obtain a lower bound for coherence cost for approximately implementing Gibbs-preserving Operations by Thermal Operations.
@@ Appendix C Construction of pairwise reversible Gibbs-preserving Operations
For a countably infinite series $\{E_{n}\}_{n}$ of real numbers, let $S_{d}(\{E_{n}\}_{n})$ be an arbitrary $d$-dimensional system equipped with Hamiltonian $H_{d}=\sum_{n=0}^{d-1}E_{n}|{n}\rangle\!\langle{n}|$. Then, for arbitrary $d\geq 3$ and $d^{\prime}\leq d$, and an arbitrary energy spectrum $\{E_{n}\}_{n}$ with $E_{i}\geq E_{j},\ \forall i,j$ such that it is not fully degenerate above the ground energy, i.e., there exists $11\leq i\leq d-$ 1 such that $E_{i+1}>E_{i}$, there exists a pairwise reversible Gibbs-preserving operation $\Lambda:S_{d}(\{E_{n}\}_{n})\to S_{d^{\prime}}(\{E_{n}\}_{n})$ with a reversible pair $\mathbb{P}$ such that ${\cal C}(\Lambda,\mathbb{P})>0$.
@@
Appendix D State transformations requiring cost-diverging Gibbs preserving Operations
(Proof of Theorem 4)
Besides the physical implementability of a given Gibbs-preserving Operation, another significant problem is whether a given state transformation $\rho\to\sigma$ that can be realized by a Gibbs-preserving Operation can physically be achieved by a finite coherence cost. Theorem 2 or 3 does not directly address this question because there are typically many quantum channels that can realize the said state transformation. To investigate the smallest coherence cost for a given state transformation, one needs to optimize over all such quantum channels, which is generally intractable.
Moreover, it can easily be seen that an arbitrary state transformation $\rho\to\sigma$ can always be realized by a finite coherence cost because if one just prepares the state $\sigma$ in the auxiliary system equipped with the same Hamiltonian as the one for the main system and swaps $\rho$ and $\sigma$, this realizes the desired state transformation with the coherence cost ${\cal F}(\sigma)$. Therefore, we need to impose additional structure to obtain nontrivial insights.
To this end, we focus on the setting where an operation implemented by a Thermal Operation with coherence assistance is also Gibbs-preserving. This is a reasonable restriction given that we are considering a state transformation that can be realized by a Gibbs-preserving Operation. Here, we employ our methods to show that some state transformations realized by Gibbs-preserving Operations can never be achieved by a Thermal Operation and finite coherence, under the restriction that the implemented opearation itself is Gibbs-preserving.
Theorem S.6 (Theorem 4 in the main text).
$\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j},$ (67)
there is a pair $(\rho,\sigma)$ of states on $S$ and $S^{\prime}$ satisfying the following two conditions:
  • $\rho\rightarrow\sigma$ is possible by a Gibbs-preserving Operation from $S$ to $S^{\prime}$.
  • There is no Gibbs-preserving Operation $\Lambda$ such that $\Lambda(\rho)=\sigma$ and ${\cal F}_{c}^{\epsilon=0}(\Lambda)$ is finite.
Proof.
Consider the following state transition
$|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}}$ (68)
where
$|\eta_{+}\rangle_{S}$ $:=\sqrt{r}|i\rangle_{S}+\sqrt{1-r}|j\rangle_{S},$ (69)
$r$ $:=\frac{\tau_{S^{\prime},i^{\prime}}-\tau_{S,j}}{\tau_{S,i}-\tau_{S,j}}.$ (70)
We first show that there exists a Gibbs-preserving Operation that realizes the transformation (68). Consider a map $\Lambda_{0}:S\to S^{\prime}$ defined by
$\Lambda_{0}(...)$ $:=\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|...]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}+\mathrm{Tr}[(1-|\eta_{+}\rangle\langle\eta_{+}|)...]\xi_{S^{\prime}},$ (71)
$\xi_{S^{\prime}}$ $:=\frac{\tau_{S^{\prime}}-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}}{1-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]}.$ (72)
$\Lambda_{0}$ is a valid quantum channel because the fact that $\tau_{S^{\prime}}\geq\left(\tau_{S^{\prime},i}-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]\right)|{i}\rangle\!\langle{i}|=0$ ensures $\xi_{S^{\prime}}\geq 0$. One can also directly see that $\Lambda_{0}$ is Gibbs-preserving satisfying $\Lambda_{0}(\tau_{S})=\tau_{S^{\prime}}$.
It now suffices to show that an arbitrary Gibbs-preserving Operation $\Lambda:S\to S^{\prime}$ that realizes the transformation (68) must come with diverging coherence cost, i.e., ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$. Let $\{M_{k}\}_{k}$ be a Kraus representation of $\Lambda$. We write each Kraus operator with respect to energy eigenbases $\{|l\rangle_{S}\}_{l=1}^{d_{S}}$ of $S$ and $\{|m\rangle_{S^{\prime}}\}_{m=1}^{d_{S^{\prime}}}$ of $S^{\prime}$ as
$M_{k}=$ $a_{k}|i^{\prime}\rangle_{S^{\prime}}\langle\eta_{+}|_{S}+\sum_{m\neq i^{\prime}}b_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{+}|_{S}+c_{k}|i^{\prime}\rangle_{S^{\prime}}\langle\eta_{-}|_{S}$
$+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{l\neq i,j}e_{k,l}|i^{\prime}\rangle_{S^{\prime}}\langle l|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S},$ (73)
where $|\eta_{-}\rangle:=\sqrt{1-r}|j\rangle_{S}-\sqrt{r}|i\rangle_{S}$.
The assumption $\Lambda(|{\eta_{+}}\rangle\!\langle{\eta_{+}}|)=|{i^{\prime}}\rangle\!\langle{i^{\prime}}|$ imposes
$b_{k,m}=0,\enskip\forall k,\enskip\forall m\mbox{ s.t. }m\neq i^{\prime}.$ (74)
Furthermore, $\sum_{k}M^{\dagger}_{k}M_{k}=\mathds{1}_{S}$ gives
$\sum_{k}|a_{k}|^{2}$ $=\sum_{k}\langle\eta_{+}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=1,$ (75)
$\sum_{k}c^{*}_{k}a_{k}$ $=\sum_{k}\langle\eta_{-}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,$ (76)
$\sum_{k}e^{*}_{k,l}a_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j$ (77)
$\sum_{k}e^{*}_{k,l}c_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{-}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (78)
Therefore, we obtain
$\tau_{S^{\prime},i^{\prime}}=$ $\langle i^{\prime}|_{S^{\prime}}\tau_{S^{\prime}}|i^{\prime}\rangle_{S^{\prime}}$
$=$ $\sum_{k}\langle i^{\prime}|_{S^{\prime}}M_{k}\tau_{S}M^{\dagger}_{k}|i^{\prime}\rangle_{S^{\prime}}$
$=$ $\sum_{k}|a_{k}|^{2}\langle\eta_{+}|\tau_{S}|\eta_{+}\rangle_{S}+\sum_{k}|c_{k}|^{2}\langle\eta_{-}|\tau_{S}|\eta_{-}\rangle_{S}+\sum_{k}\sum_{l\neq i,j}|e_{k,l}|^{2}\langle l|_{S}\tau_{S}|l\rangle_{S}$
$+2\mathrm{Re}\left(\sum_{k}c^{*}_{k}a_{k}\langle\eta_{+}|\tau_{S}|\eta_{-}\rangle\right)+2\mathrm{Re}\left(\sum_{l\neq i,j}\sum_{k}e^{*}_{k,l}a_{k}\langle\eta_{+}|\tau_{S}|l\rangle\right)+2\mathrm{Re}\left(\sum_{l\neq i,j}\sum_{k}e^{*}_{k,l}c_{k}\langle\eta_{-}|\tau_{S}|l\rangle\right)$
$=$ $r\tau_{S,i}+(1-r)\tau_{S,j}+\sum_{k}|c_{k}|^{2}((1-r)\tau_{S,i}+r\tau_{S,j})+\sum_{k}\sum_{l\neq i,j}|e_{k,l}|^{2}\tau_{S,l}$
$=$ $\tau_{S^{\prime},i^{\prime}}+\sum_{k}|c_{k}|^{2}((1-r)\tau_{S,i}+r\tau_{S,j})+\sum_{k}\sum_{l\neq i,j}|e_{k,l}|^{2}\tau_{S,l}.$ (79)
Because of $(1-r)\tau_{S,i}+r\tau_{S,j}>0$ and $\tau_{S,l}>0$, we get
$c_{k}$ $=0,\enskip\forall k$ (80)
$e_{k,l}$ $=0,\enskip\forall k,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (81)
Therefore, we can rewrite $M_{k}$ as
$M_{k}=$ $a_{k}|i^{\prime}\rangle_{S^{\prime}}\langle\eta_{+}|_{S}+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S}.$ (82)
This particularly implies that
$\langle i^{\prime}|_{S^{\prime}}\Lambda(\psi_{S})|i^{\prime}\rangle_{S^{\prime}}=0,\enskip\forall|\psi\rangle_{S}\mbox{ s.t. }\langle{\eta_{+}}|{\psi}\rangle_{S}=0.$ (83)
We employ this to construct a reversible input state pair that shows diverging coherence cost via Theorem 2. Let =$\mathbb{P}=\{|{\eta_{+}}\rangle\!\langle{\eta_{+}}|,|{\eta_{-}}\rangle\!\langle{\eta_{-}}|\}$. Then, one can see that $\mathbb{P}$ is a reversible pair by considering a recovery channel
${\cal R}(...)=\mathrm{Tr}[|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|...]|\eta_{+}\rangle\langle\eta_{+}|_{S}+\mathrm{Tr}[(1-|i^{\prime}_{S^{\prime}}\rangle\langle i^{\prime}_{S^{\prime}}|)...]|\eta_{-}\rangle\langle\eta_{-}|_{S},$ (84)
which perfectly recovers $|\eta_{\pm}\rangle$ because of (83). We can also directly check that ${\cal C}(\Lambda,\mathbb{P})>0$ as follows. Note that the operator $\Lambda^{\dagger}(H^{\prime})$ satisfies
$\langle\eta_{+}|_{S}\Lambda^{\dagger}(H^{\prime})|\eta_{-}\rangle_{S}$ $=\sum_{k}\langle\eta_{+}|_{S}M^{\dagger}_{k}H^{\prime}M_{k}|\eta_{-}\rangle_{S}$
$=\sum_{m\neq i^{\prime}}\sum_{k}a^{*}_{k}E_{S^{\prime},m}d_{k,m}\delta_{i^{\prime}m}$
$=0.$ (85)
Therefore,
${\cal C}(\Lambda,\mathbb{P})^{2}$ $=|\langle\eta_{+}|H|\eta_{-}\rangle_{S}|^{2}$
$=|\langle\eta_{+}|(E_{S,i}|i\rangle\langle i|_{S}+E_{S,j}|j\rangle\langle j|_{S})|\eta_{-}\rangle|^{2}$
$=r(1-r)\left(E_{S,i}-E_{S,j}\right)^{2}$
$>0$ (86)
where the last inequality is because of the assumption $\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j}$, which ensures $E_{S,i}-E_{S,j}\neq 0$ and $10<r<$ 1. Applying Theorem 2 to $\Lambda$ and $\mathbb{P}$ results in ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$, concluding the proof. ∎
Appendix E Upper bound for the coherence cost of Eq. (2)